OpenStudy (anonymous):

The volume of a sphere is 3,000π m3. What is the surface area of the sphere to the nearest square meter?

OpenStudy (lgbasallote):

volume = 4/3 pi r ^3 \[3,000 \pi = \frac 43 \pi r^3\] \[9000\pi = 4\pi r^3\] \[2250 \pi = \pi r^3\] \[2250 = r^3\] \[r = 47.43\] surface area = 4pi r^2 \[SA = 4 \pi (47.43)^2\] got it?

OpenStudy (anonymous):

No lol, I got 28255.037544

OpenStudy (lgbasallote):

oh lol sorry...i took the sqrt of 2250 it should be cube root

OpenStudy (lgbasallote):

radius should be 13.1

OpenStudy (lgbasallote):

so surface area = 4 pi (13.1)^2

OpenStudy (lgbasallote):

yup that's a better answer

OpenStudy (lgbasallote):

do you get it?

OpenStudy (anonymous):

Yessir thankie

OpenStudy (lgbasallote):

<tips hat>